OpenStudy (lgbasallote):

\[\mathcal L ^{-1} \lbrace \frac{s}{s^2 + 4s + 4} \rbrace\]

OpenStudy (lgbasallote):

i got as far as \[\implies \mathcal L ^{-1} \left \lbrace \frac{s}{(s+2)^2} \right \rbrace\] but after that i got nothing

OpenStudy (lgbasallote):

maybe +2 - 2 in the numerator?

OpenStudy (lgbasallote):

okay...i got \[e^{-2t} \cos t - 2e^{-2t} \sin t\] that right?

OpenStudy (experimentx):

ask wolf

OpenStudy (lgbasallote):

i wish i could..idk how to do laplace transforms in wolf

OpenStudy (anonymous):

ok let me check it.

OpenStudy (experimentx):